Trace the curve y= 8/4-x^2 , and state all the properties you use to trace it.
1
Expert's answer
2021-10-28T11:07:13-0400
"y=\dfrac{8}{4-x^2}"


1. "4-x^2\not=0=>x\not=\pm2"

Domain: 

"(-\infin;-2)\cup(-2,2)\cup(2, \infin)"


2. "y-" intersection:

"x=0, y=\dfrac{8}{4-(0)^2}=2"

Point "(0, 2)"

"x-" intersection(s):

"y=0, \dfrac{8}{4-x^2}=0," No solution

There are no "x" -intersections.


3. "y(-x)=\dfrac{8}{4-(-x)^2}=\dfrac{8}{4-x^2}=y(x)"

The function "y=\dfrac{8}{4-x^2}" is even on its domain. The garph is symmetric with respect to "y" -axis.


4.


"\lim\limits_{x\to -2^-}\dfrac{8}{4-x^2}=-\infin"

"\lim\limits_{x\to -2^+}\dfrac{8}{4-x^2}=\infin"

"\lim\limits_{x\to 2^-}\dfrac{8}{4-x^2}=\infin"

"\lim\limits_{x\to 2^+}\dfrac{8}{4-x^2}=-\infin"

Vertical asymptotes: "x=-2, x=2."



"\lim\limits_{x\to -\infin}\dfrac{8}{4-x^2}=0"

"\lim\limits_{x\to \infin}\dfrac{8}{4-x^2}=0"

Horizontal asymptote: "y=0."


5.

"y'=(\dfrac{8}{4-x^2})'=\dfrac{16x}{(4-x^2)^2}"

Find the critical number(s):


"y'=0=>\dfrac{16x}{(4-x^2)^2}=0=>x=0, x\not=\pm2"

If "x<-2, y'<0, y" decreases.

If "-2<x<0, y'<0, y" decreases.

If "0<x<2, y'>0, y" increases.

If "x>2, y'>0, y" increases.


"y(0)=\dfrac{8}{4-(0)^2}=2"

The function "y" has a local minimum with value of "2" at "x=0."


6.

"y''=(\dfrac{16x}{(4-x^2)^2})'=\dfrac{16((4-x^2)^2-2x(4-x^2)(-2x))}{(4-x^2)^4}"

"=\dfrac{16((4-x^2)^2-2x(4-x^2)(-2x))}{(4-x^2)^4}"

"=\dfrac{16(4-x^2+4x^2)}{(4-x^2)^3}=\dfrac{16(4+3x^2)}{(4-x^2)^3}"


If "x<-2, y'<0, y" is concave down.

If "-2<x<2, y''>0, y" is concave up.

If "x>2, y''<0, y" is concave down.


7. Sketch the graph.